Calculate the minimum deviation produced by a 60° glass prism if the refractive index of the glass is

    1.50. What is the angle of incidence at which minimum deviation occurs?



1
Expert's answer
2021-03-08T08:32:42-0500

Answer

Angle of incidence is given by using snell law

"I=\sin^{-1}({\mu sin R})"

"=\sin^{-1}({1.5sin 60°})"

=49°


Now minimum deviation

"\delta_m=2I-A
 \\=2\times49°-60°\\=38°"